OpenStudy (anonymous):

Implicit differentiation involving trig and finding equation of tangent line. Please help

OpenStudy (anonymous):

\[ysin(2x) = xcos(2y)\] . find equation of tangent line at the point \[\frac{ \pi }{ 2 }, \frac{ \pi }{ 4 }\]

hartnn (hartnn):

so, if you find dy/dx and then put x = pi/2, y= pi/4 you will get slope of tangent at that point. now you have slope and a point on a line, it would be easy to find its equation.

OpenStudy (anonymous):

i solved for dy/dx. unsure of how to proceed from there to get the gradient of my equation of the tangent line.

hartnn (hartnn):

put x = pi/2, y= pi/4

OpenStudy (anonymous):

yeah i know that, how would i solve for pi/2 and pi/4 for example cos(pi/2) ?

hartnn (hartnn):

cos (pi/2) = 0

OpenStudy (anonymous):

here is my dy/dx: \[\frac{ dy }{ dx } = \frac{ \cos(2y) - 2y.\cos(2x) }{ \sin(2x) + 2x.\sin(2y) }\]

hartnn (hartnn):

and its absolutely correct :)

OpenStudy (anonymous):

so i would plug in for x and y respectivley with the point given?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

what if it doesnt factor out nicely would I still use it as my m for y=mx+c

OpenStudy (anonymous):

going to solve for x,y now, gimme 2min please :)

hartnn (hartnn):

sure, take your time :)

hartnn (hartnn):

\(\huge m = \frac{ \cos(\pi/2) - 2(\pi/4).\cos(\pi) }{ \sin(\pi) + 2(\pi/2).\sin(\pi/2) }\)

OpenStudy (anonymous):

slight mistake in your cos(pi/2). Should be cos(2.pi/2) which can be simplified to cos(pi). correct?

hartnn (hartnn):

y=pi/4

OpenStudy (anonymous):

plugged in for points pi/2 and pi/4. Then simplified further. Then used trig to find the arc of the sin/cos to get a number. \[\frac{ \cos(\pi)-\frac{ \pi }{ 2 }\cos(\pi) }{ \sin(\pi) + \pi.\sin(\frac{ \pi }{ 2 }) }\]

hartnn (hartnn):

y = pi/4 cos 2y = cos 2*pi/4 = cos pi/2

OpenStudy (anonymous):

ahh yes you are correct.

hartnn (hartnn):

so, m=.. ?

OpenStudy (anonymous):

\[\frac{ \cos(\ \frac{ \pi }{ 2 })-\frac{ \pi }{ 2 }\cos(\pi) }{ \sin(\pi) + \pi.\sin(\frac{ \pi }{ 2 }) }\]

hartnn (hartnn):

simplify that, can u ?

OpenStudy (anonymous):

solved further using trig: \[\frac{ -1-\frac{ \pi }{ 2 } (-1)}{ 0+\pi(1)}\]

OpenStudy (anonymous):

please double check that for me

hartnn (hartnn):

cos pi/2 = 0

hartnn (hartnn):

everything else is correct

OpenStudy (anonymous):

ah yes, forgot to edit my initial calculation

OpenStudy (anonymous):

m = 1/2

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

ahhh. now I simply use y = mx + c. y = 0.5(x)+c plugin points pi/2, pi/4 for x and y respectivley: pi/4 = 0.5(pi/2) + c c = 0.5

OpenStudy (anonymous):

\[\therefore y=\frac{ 1 }{ 2 }x\]

hartnn (hartnn):

you could also use y- y1 = m (x-x1) y-pi/4 = 0.5 (x- pi/2) y- pi/4 = 0.5 x - pi/4 y= x/2 yes, its correct :)

OpenStudy (anonymous):

thanks so much once again!

hartnn (hartnn):

welcome once again ^_^